Find the quotient and remainder when : x^4-2x2+5x-4 is divided by x+3. Use either long or sythetic division
Answer by venugopalramana(3286)

Find the quotient and remainder when : x^4-2x2+5x-4 is divided by x+3. Use either long or sythetic division
X+3=0....X=-3
-3...|..1.....+0.......-2.......+5........-4
.....|..0.....-3........9.......-21.......48
------------------------------------------------------
........1.....-3........7.......-16.......44
REMAINDER=44
QUOTIENT=X^3-3X^2+7X-16